40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 7 14:39:49 2013, Jenne, Update, LSC, Lock Acquisition Game Plan LockAcquisition_7Nov2013.pdf
    Reply  Fri Nov 8 18:12:21 2013, Jenne, Update, LSC, PRFPMI: Not crossing any resonances 2f_resonances_Xarm_CARMsweep.pngAllModFreq_resonances_Xarm_CARMsweep.png
    Reply  Tue Nov 12 16:49:22 2013, Jenne, Update, LSC, Xend QPD and Whitening board pulled IMG_1262.JPGIMG_1260.JPGIMG_1264.JPGIMG_1276.JPG
       Reply  Wed Nov 20 17:05:15 2013, Jenne, Update, LSC, Xend QPD and Whitening board replaced 
       Reply  Wed Dec 18 00:26:15 2013, Jenne, Update, LSC, Xend QPD schematic investigation 
          Reply  Wed Dec 18 11:32:34 2013, rana, Update, LSC, Xend QPD schematic investigation 
    Reply  Wed Nov 13 01:35:40 2013, Jenne, Update, LSC, PRM motion causing trouble? xarmRIN_vs_CARM.pngASportOverlap_moveGaussianBeams.png
       Reply  Wed Nov 13 09:31:15 2013, Gabriele, Update, LSC, PRM motion causing trouble? 
          Reply  Wed Nov 13 19:22:58 2013, Jenne, Update, LSC, PRM motion correlated to intracavity power POPDC_sensors_30oct2013.pngAOMmodulated_POPDC_OSEM.pdf
             Reply  Thu Nov 14 02:50:43 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OL_TF_model.pngPRM_OLPIT_TF.pdfPRM_OLYAW_TF.pdf
                Reply  Mon Nov 18 21:02:54 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OLG_NewOld.pngPRM_oplev_improvement.pdf
                   Reply  Mon Nov 18 21:40:24 2013, Koji, Update, LSC, PRM oplev measured and modeled TF 
                      Reply  Tue Nov 19 14:47:44 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF 
             Reply  Thu Nov 14 02:55:26 2013, rana, Update, SUS, PRM motion correlated to intracavity power 
                Reply  Thu Nov 14 11:41:19 2013, Steve, Update, SUS, PRM sensors effected by IR IRon_off.png
                   Reply  Thu Nov 14 14:35:12 2013, Steve, Update, SUS, IR effect on MC and PRM sensors MC_PRM_IReffect.png
                      Reply  Fri Dec 6 10:03:07 2013, Steve, Update, SUS, IR effect on MC sensors only 10minPMCnotLocked.png6Dec2013.png
Message ID: 9378     Entry time: Wed Nov 13 19:22:58 2013     In reply to: 9373     Reply to this: 9382   9383
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRM motion correlated to intracavity power 

[Gabriele, Jenne]

Nic and Evan put the ISS together (elog 9376), and we used an injection into the error point (?) to modulate the laser power before the PMC (The AOM had a bias offset, but there is no loop).  This gives us some RIN, that we can try to correlate with the PRM OSEM sensors. 

We injected several lines, around 100, 200, 500 and 800 Hz.  For 100, 200 and 800 Hz lines, we see a ratio between POPDC and the OSEM sensors of 1e-4, but at 500 Hz, the ratio was more like 1e-3. We're not sure why this ratio difference exists, but it does.  These ratios were true for the 4 face OSEMs.  The side OSEM saw a slightly smaller signal.  

For these measurements, the PRMI was sideband locked, and we were driving the AOM with an amplitude of 10,000 counts (I don't know what the calibration is between counts and actual drive, which is why we're looking at the POPDC to sensor *ratio*).  

To get a more precise number, we may want to consider locking the PRMI on carrier, so we have more power in the cavity, and so more signal in the OSEMs.  

These ratios look, by eye, similar to the ratios we see from the time back on 30 Oct when we were doing the PRMI+2arms test, and the arms were resonating about 50 units.  So, that is nice to see some consistency.

AOMmodulated_POPDC_OSEM.pdf

This time series is from 1067163395 + 27 seconds, from 30 Oct 2013 when we did the PRMI+2arms.

POPDC_sensors_30oct2013.png

 


Ideas to go forward:

We should think about chopping the OSEM LEDs, and demodulating the PD sensors. 

We should also take a look in the chamber with a camera from the viewport on the north side of the BS chamber, to see if we see any flashes in the chamber that could be going into the OSEMs, to see where we should maybe put aluminum foil shields.

ELOG V3.1.3-